# references.bib

### v1

@@ Abstract
Language-image pre-training has demonstrated strong performance in 2D medical imaging, but its success in 3D modalities such as CT and MRI remains limited due to the high computational demands of volumetric data, which pose a significant barrier to training on large-scale, uncurated clinical studies. In this study, we introduce Hierarchical attention for Language-Image Pre-training (HLIP), a scalable pre-training framework for 3D medical imaging. HLIP adopts a lightweight hierarchical attention mechanism inspired by the natural hierarchy of radiology data: slice, scan, and study. This mechanism exhibits strong generalizability, e.g., +4.3% macro AUC on the Rad-ChestCT benchmark when pre-trained on CT-RATE. Moreover, the computational efficiency of HLIP enables direct training on uncurated datasets. Trained on 220K patients with 3.13 million scans for brain MRI and 240K patients with 1.44 million scans for head CT, HLIP achieves state-of-the-art performance, e.g., +32.4% balanced ACC on the proposed publicly available brain MRI benchmark Pub-Brain-5; +1.4% and +6.9% macro AUC on head CT benchmarks RSNA and CQ500, respectively. These results demonstrate that, with HLIP, directly pre-training on uncurated clinical datasets is a scalable and effective direction for language-image pre-training in 3D medical imaging. The code is available at https://github.com/Zch0414/hlip
@@ 1 Introduction
Language-supervised pre-training is well-suited for radiology, where each study comprises medical images paired with a corresponding radiologist’s report. This natural alignment between visual and textual information has motivated the adaptation of language-image pre-training methods, such as CLIP [zhang2022contrastive, radford2021learning], to learn clinically meaningful radiology representations. CLIP-based models are especially notable for their clinical utility: they demonstrate strong zero-shot transfer performance on diagnostic tasks [hamamci2024developing, blankemeier2024merlin], and their encoders consistently improve performance on multimodal learning benchmarks [bai2024m3d, shui2025large]. In the domain of 2D medical imaging, such as chest X-rays, language-supervised pre-training is a key driver for integrating computer vision into clinical workflows [boecking2022making, wang2022medclip, tiu2022expert].
@@
However, language-supervised pre-training in 3D medical imaging has yet to reach the scale or performance demonstrated in 2D modalities. For instance, chest X-ray CLIP models have been trained on over 500K 2D images [wang2022medclip, tiu2022expert, you2023cxr], achieving human-level performance on multiple diagnostic tasks [tiu2022expert]. In contrast, progress in 3D medical imaging remains limited, with existing models underperforming relative to their 2D counterparts [hamamci2024developing, bai2024m3d]. We attribute this performance gap between 2D and 3D language-image pre-training to three primary factors: the need for data curation and annotation, limitations in model architectures, and the complexity of 3D medical imaging studies.
@@
Computed tomography (CT) and magnetic resonance imaging (MRI) generate 3D volumetric images across various anatomical regions, including the brain, chest, and abdomen. Unlike 2D imaging, a single clinical CT or MRI study may comprise multiple 3D volumes, often called scans or sequences, each capturing different imaging protocols or orientations. As illustrated in Figure 1, a standard MRI study typically includes several sequences (e.g., T1-weighted, T2-weighted, and FLAIR), each contributing distinct diagnostic information. Similarly, CT studies often include scans acquired with varying orientations or scanner settings within the same study. Naively encoding such uncurated studies with standard visual encoders, such as the Vision Transformer (ViT) [dosovitskiy2020image], can produce on the order of $10^{4}$ tokens per study, resulting in substantial computational overhead. The sheer size and structural complexity of uncurated studies therefore present a significant barrier to scalable language-image pre-training. A common strategy is to curate datasets by having radiologists manually select a representative scan or 2D slice from each study, as shown in Figure 1 (a)  [blankemeier2024merlin, hamamci2024developing, bai2024m3d, shui2025large]. Alternatively, specialized model architectures, which depend heavily on preprocessing, have been proposed to reduce complexity[hamamci2024developing]. Yet, these bespoke designs tend to generalize and scale poorly in real-world settings and are incompatible with established ViT training protocols [radford2021learning, he2022masked, cherti2023reproducible]
Figure 1: Illustration of (a) an uncurated study for a patient. While previous work has relied on annotation and curation, HLIP enables language-image pre-training directly on such data. (b) HLIP achieves higher performance in zero-shot brain MRI disease classification while using less memory than the original ViT, which can only be trained with gradient checkpointing.
@@
In this work, we address the key barriers to scaling language-image pre-training for 3D medical imaging by directly leveraging uncurated studies and the original ViT. To effectively extract features from uncurated studies, we introduce Hierarchical attention for Language-Image Pre-training (HLIP), a lightweight adaptation that leverages the natural hierarchy of radiology data: slice, scan, and study. Unlike architectural designs such as Swin [liu2021swin], MViT [fan2021multiscale, li2022mvitv2], and Hiera [ryali2023hiera], HLIP utilizes the radiology data structure to define the attention scope. Furthermore, compared to efficiency-oriented modules such as window attention [li2022exploring], HLIP does not trade accuracy for reduced compute. As illustrated in Figure 1 (b), our method achieves higher performance while using less memory than ViT. Moreover, HLIP is orthogonal to flash attention [dao2022flashattention] and patch dropout [li2023scaling], further alleviating the computational burden associated with uncurated studies that contain multiple 3D scans.
@@
When trained on the public CT-RATE [hamamci2024developing] dataset, HLIP demonstrates strong generalizability, outperforming the state-of-the-art model [liu2023large] by 4.3% macro AUC on the external Rad-ChestCT [draelos2021machine] dataset. Furthermore, HLIP enables scalable and effective vision-language pre-training on health system-scale datasets. Specifically, HLIP trained on our health system outperforms 2D foundation models [zhang2023biomedclip, nie2025conceptclip] on the proposed publicly available brain MRI benchmark, Pub-Brain-5, by 32.4% balanced ACC; and surpasses the head CT foundation model [zhu20253d] by 1.4% and 6.9% macro AUC on the RSNA [flanders2020construction] and CQ500 [chilamkurthy2018deep] benchmarks, respectively. Our key contributions lie in three-fold:
@@
  • We introduce HLIP, a hierarchical attention mechanism that leverages the natural structure of radiology data to enable effective and scalable language-image pre-training for 3D medical imaging.
@@
  • We conduct the largest-scale training for 3D medical imaging to date, using 220K patients with 3.13 million scans for brain MRI and 240K patients with 1.44 million scans for head CT.
@@
  • We demonstrate state-of-the-art performance on multiple benchmarks spanning diverse modalities and anatomical regions, including chest CT, head CT, and brain MRI.
@@ 2 Related Work
Language-Image Pre-training in Radiology has been developed for 2D and 3D medical imaging, facilitated by public datasets such as CheXpert [irvin2019chexpert], MIMIC-CXR [johnson2019mimic], and CT-RATE [hamamci2024developing]. In 2D imaging, techniques such as local alignment [huang2021gloria, wang2022multi, muller2022joint], knowledge enhancement [wang2022medclip, wu2023medklip], and longitudinal analysis [bannur2023learning] have been explored. CheXZero [tiu2022expert] demonstrates strong empirical results comparable to those of human experts in the domain of chest X-rays. BiomedCLIP [zhang2023biomedclip] and ConceptCLIP [nie2025conceptclip] have been trained on more than 15 million sample pairs, achieving strong performance across modalities, including radiology. In addition, other works [boecking2022making, you2023cxr, zhang2023knowledge] have also contributed significant insights and achieved impressive performance.
3D medical imaging offers a more comprehensive view of anatomical structures. However, due to its computational cost and limited data availability, several studies [cao2024bootstrapping, wang2024enhancing, he2024unified] focus on bridging the domain gap between 2D and 3D. For example, UniMedI [he2024unified] learns a shared feature space for both 2D and 3D modalities and demonstrates improvements across both. BIUD [cao2024bootstrapping] distills 3D representations from a well-trained 2D model [tiu2022expert], significantly improving data efficiency.
@@
More recently, aided by advances in hardware and the availability of public datasets [hamamci2024developing], several studies [hamamci2024developing, bai2024m3d, shui2025large, lai2025bridged] have performed language-image pre-training on 3D imaging, surpassing methods [cao2024bootstrapping] that still rely on 2D representations. Specifically, CT-CLIP [hamamci2024developing] has been pre-trained on 47,149 paired chest CT scans and reports. M3D [bai2024m3d] explores a versatile multi-modal large language model designed for universal 3D medical imaging analysis. fVLM [shui2025large] proposes a fine-grained pre-training framework that relies on segmentation to perform organ alignment. However, while these methods have demonstrated promising performance on various tasks such as zero-shot abnormality detection and report generation, several factors limit their training scalability and real-world applicability. For example, CT-CLIP and M3D are constrained to radiology studies containing only a single imaging scan. Such studies rely heavily on human annotation and also fail to reflect real-world scenarios, where studies typically include multiple scans. fVLM further depends on a segmentation model, which introduces bias from segmentation quality and largely limits its scalability and generalizability. Moreover, these methods suffer from inefficient modeling of 3D medical imaging, leading to small batch sizes (e.g., 48), which are insufficient for effective language-image pre-training [radford2021learning].
@@
Efficient Language-Image Pre-training is crucial, as it directly affects the number of sample pairs seen during training and contrasted per batch, two key factors of model capacity. Merlin [blankemeier2024merlin] and CT-CLIP [hamamci2024developing] adopt hierarchical architectures; however, as analyzed in Appendix E.2, these models [liu2021swin, hamamci2024developing] may be less efficient than commonly assumed when compared to the original ViT. From the architecture perspective, components such as relative position embedding are expensive for 3D inputs and is not compatible with recent advancements like flash attention [dao2022flashattention]. From the training perspective, FLIP [li2023scaling] demonstrates a favorable trade-off by randomly removing 50% of tokens during training. However, models that require a fixed activation shape [liu2021swin, li2022mvitv2, li2022exploring, ryali2023hiera] cannot benefit from this strategy. Moreover, many works in radiology [matsoukas2022makes, tiu2022expert, zhao2024part, shui2025large] have demonstrated the benefits of the universal feature learned by MAE [he2022masked]. Therefore, exploring an effective approach to modeling 3D medical imaging with the original ViT [dosovitskiy2020image] is a promising research direction and stands to benefit from recent advancements such as flash attention, patch dropout, and pre-trained models.
@@ 3 Method
Our goal is to perform language-image pre-training on uncurated studies using the original ViT [dosovitskiy2020image]. Given a study $S\in\mathbb{R}^{M\times 1\times D\times H\times W}$, where $M$ denotes the number of single-channel 3D scans, and $D$, $H$, and $W$ represent the depth, height, and width dimensions of each scan, respectively. Following the ViT [dosovitskiy2020image], each scan is divided into a grid of non-overlapping volumes of size $(\frac{D}{d},\frac{H}{h},\frac{W}{w})$. These volumes are projected into visual tokens $F_{s}\in\mathbb{R}^{N\times c}$, where $N=M\times d\times h\times w$ is the total number of tokens and $c$ denotes the number of channels. As $N$ can be on the order of $10^{4}$, computing self-attention over all tokens throughout the ViT backbone is prohibitive in memory. We introduce a hierarchical attention mechanism guided by the inherent data hierarchy of radiology studies, enabling more effective self-attention while preserving the ViT architecture. The strength of our approach lies in enabling scalable language-image pre-training on uncurated studies while leveraging the established training protocol of ViT [radford2021learning, he2022masked, cherti2023reproducible].
Figure 2: Illustration of (a) the radiology data hierarchy for a single patient, including the study, single scan, and adjacent slices. Our hierarchical attention mechanism mirrors this hierarchy and computes self-attention independently within each level. (b) Our HLIP framework incorporates a visual encoder that performs attention at different levels. In practice, lightweight slice or scan attention with a few study attention layers suffices to extract features from the full radiology study.
@@ 3.1 Hierarchical Attention Mechanism
The radiology study exhibits an inherent hierarchical structure for each patient, and our proposed hierarchical attention mechanism mirrors this structure. As illustrated in Figure 2, the study for an individual patient comprises three levels:
@@
  • The scan contains $D$ slices. While a single scan conveys only partial context of the full radiology study, it still captures the complete extent of the target pathology.
@@
  • Study attention computes a single self-attention operation over all $N$ tokens in a study. The I/O complexity [aggarwal1988input] of the study attention is $\Omega(N^{2}+N\times c)$ [dao2022flashattention], as detailed in Appendix E.
@@
This contrasts with existing methods that modify attention computation using convolution [fan2021multiscale, li2022mvitv2], pooling [ryali2023hiera], or shifted windows [liu2021swin], all of which require either regular activation shapes or attention masks. Our mechanism relies solely on a simple reshape operation, as there is no overlap between different scans or groups of adjacent slices [^1] With the original attention layer unchanged, study attention takes the input of size $(B,\ M\times d\times h\times w,\ c)$; scan attention takes the input of size $(B\times M,\ d\times h\times w,\ c)$; and slice attention takes the input of size $(B\times M\times d,\ h\times w,\ c)$, where $B$ denotes the batch size., allowing greater flexibility for radiology studies that comprise multiple scans. As shown in Figure 2, each layer in the ViT can perform attention at any level. In practice, we evenly divide the ViT backbone into four subsets of layers (e.g., three layers per subset for the 12-layer ViT-B) and apply study attention only to the last layer of each subset, while the remaining layers perform the lighter scan or slice attention.
HLIP integrates seamlessly with recent efficiency advances like patch dropout [li2023scaling] and flash attention [dao2022flashattention]. Furthermore, unlike window attention [li2022exploring], our scan and slice attentions do not trade performance for efficiency as they preserve all diagnostic features.

### v2

# Towards Scalable Language-Image Pre-training for 3D Medical Imaging

@@ Abstract
The scalability of current language-image pre-training for 3D medical imaging, such as CT and MRI, is constrained by the need for radiologists to manually curate raw clinical studies. In this work, we pioneer pre-training directly on uncurated studies, which both aligns more closely with the radiologist’s workflow and provides a natural path to scalability. However, the unique structure of such data presents new challenges for existing model architectures, which were originally designed for 2D slices or single 3D scans. To address this, we introduce a novel hierarchical attention mechanism inspired by the intrinsic hierarchy of radiology data: slice, scan, and study. We denote our framework as Hierarchical attention for Language-Image Pre-training (HLIP). Trained on 220K studies with 3.13 million scans for brain MRI and 240K studies with 1.44 million scans for head CT, HLIP achieves state-of-the-art performance, e.g., +10.5% balanced ACC on the proposed publicly available brain MRI benchmark Pub-Brain-5; +8.3% and +1.7% macro AUC on head CT benchmarks CQ500 and RSNA, respectively. HLIP also exhibits strong generalizability on existing 3D medical language-image pre-training benchmarks, e.g., +4.3% macro AUC on the Rad-ChestCT benchmark when pre-trained on CT-RATE. These results demonstrate that, with HLIP, directly pre-training on uncurated clinical datasets is a scalable and effective direction for language-image pre-training in 3D medical imaging. The code is available at https://github.com/Zch0414/hlip
@@ 1 Introduction
Language-supervised pre-training is well-suited for radiology, where each study comprises medical images paired with a corresponding radiologist’s report. This natural alignment between visual and textual information has motivated the adaptation of language-image pre-training methods, such as CLIP [1, 2], to learn clinically meaningful radiology representations. CLIP-based models are especially notable for their clinical utility: they demonstrate strong zero-shot transfer performance on diagnostic tasks [3, 4], and their encoders consistently improve performance on multimodal learning benchmarks [5, 6]. In the domain of 2D medical imaging, such as chest X-rays, language-supervised pre-training is a key driver for integrating computer vision into clinical workflows [7, 8, 9].
@@
However, language-image pre-training in 3D medical imaging has yet to reach the scale or performance demonstrated in 2D modalities. For instance, chest X-ray CLIP models have been trained on a 500K corpus [8, 9, 10], achieving human-level performance on multiple diagnostic tasks [9]. In contrast, progress in 3D medical imaging remains limited, with existing models underperforming relative to their 2D counterparts [3, 5]. We attribute the performance gap between 2D and 3D language-image pre-training to two primary factors: data annotation that constrains the training scale, and architectural limitations arising from the complexity of 3D medical imaging.
@@
Computed tomography (CT) and magnetic resonance imaging (MRI) generate 3D volumetric images across various anatomical regions, including the brain, chest, and abdomen. As illustrated in Figure 1, a standard MRI study typically includes several sequences (e.g., T1-weighted, T2-weighted, and FLAIR), each contributing distinct diagnostic information. Similarly, CT studies often include scans acquired with varying orientations or scanner settings within the same study. To perform language-image pre-training for such data, a common strategy is to curate datasets by having radiologists manually select a representative scan or slice from each study, as shown in Figure 1(a) [4, 3, 5, 6], which presents a significant barrier to the scalability. In contrast, pre-training on uncurated studies aligns more closely with real-world practice and readily expands the data scale, as it imposes no additional burden on radiologists. Despite this, the unique structure of such data presents new challenges for current visual encoders, which were originally designed for 2D images or single 3D scans [11, 12, 13]. As shown in Figure 1(b), even trained on a large-scale domain-specific dataset, naively encoding uncurated studies with the Vision Transformer (ViT) [11] results in only comparable zero-shot transferability to the state-of-the-art (SOTA) biomedical CLIP [14] trained on the PubMed corpus. Particularly, encoding the entire study can produce tokens on the order of $10^{4}$, which both incurs substantial computational overhead and limits performance [15].
Figure 1: Illustration of (a) an uncurated study for a patient. While previous work has relied on annotation and curation, HLIP enables language-image pre-training directly on uncurated data. (b) Despite training on large-scale domain-specific datasets, naively modeling the uncurated study with a vanilla ViT, e.g., by randomly selecting a scan at each training step, encoding scans independently before study aggregation, or directly encoding the entire study, yields performance only comparable to the SOTA trained on PubMed corpus, whereas HLIP outperforms these by a large margin.
@@
In this work, we first address the key barriers to scaling language-image pre-training for 3D medical imaging by pioneering the use of uncurated studies. Second, to effectively extract features from such data, we introduce a novel hierarchical attention mechanism leverages the natural hierarchy of radiology data: slice, scan, and study. We name this framework Hierarchical attention for Language-Image Pre-training (HLIP). Unlike architectural designs such as Swin [12], MViT [16, 17], and Hiera [13], HLIP leverages the inherent data structure to define the attention scope, making it suitable for uncurated studies that contain multiple 3D scans. Compared to window attention that captures only local features, slice or scan attention can capture all diagnostic features while also providing constructive priors for learning study representations. Moreover, slice and scan attention is already much lighter than study attention, and such minimal adaptation of the original ViT remains orthogonal to flash attention [18] and patch dropout [19], further reducing the computational burden.
@@
Trained on our health system, HLIP outperforms the SOTA [14] on the proposed publicly available brain MRI benchmark, Pub-Brain-5, by 10.5% balanced ACC; and surpasses the head CT foundation model [20] by 8.3% and 1.7% macro AUC on the CQ500 [21] and RSNA [22] benchmarks, respectively. HLIP also demonstrates strong generalizability on the curated 3D medical language-image pre-training benchmark CT-RATE [3], which contains only one scan per study, outperforming the SOTA [23] by 4.3% macro AUC on the external evaluation Rad-ChestCT [24].
Our paper makes the following contributions:
@@
  • We introduce HLIP, an effective and scalable language-image pre-training framework for uncurated 3D medical imaging, that leverages a novel hierarchical attention mechanism derived from the natural structure of radiology data.
@@
  • We conduct the largest-scale training for 3D medical imaging to date, using 220K studies with 3.13 million scans for brain MRI and 240K studies with 1.44 million scans for head CT.
@@
  • We demonstrate the state-of-the-art performance on multiple benchmarks spanning diverse modalities and anatomical regions, including brain MRI, head CT and chest CT.
  • We release the following assets to the public: a brain MRI benchmark for zero-shot classification, an effective language-image pre-training implementation for 3D medical imaging, the pre-training recipe, and model checkpoints.
@@ 2 Related Work
Language-Image Pre-training in Radiology has been developed for 2D and 3D medical imaging, facilitated by public datasets such as CheXpert [25], MIMIC-CXR [26], and CT-RATE [3]. In 2D imaging, techniques such as local alignment [27, 28, 29], knowledge enhancement [8, 30], and longitudinal analysis [31] have been explored. CheXZero [9] demonstrates strong empirical results comparable to those of human experts in the domain of chest X-rays. BiomedCLIP [32] and ConceptCLIP [14] have been trained on more than 15 million sample pairs, achieving strong performance across modalities, including radiology. In addition, other works [7, 10, 33] have also contributed significant insights and achieved impressive performance. 3D medical imaging offers a more comprehensive view of anatomical structures. However, due to its computational cost and limited data availability, several studies [34, 35, 36] focus on bridging the domain gap between 2D and 3D. For example, UniMedI [36] learns a shared feature space for both 2D and 3D modalities and demonstrates improvements across both. BIUD [34] distills 3D representations from a well-trained 2D model [9], significantly improving data efficiency.
@@
More recently, aided by advances in hardware and the availability of public datasets [3], several studies [3, 5, 6, 37] have performed language-image pre-training on 3D imaging, surpassing methods [34] that still rely on 2D representations. Specifically, CT-CLIP [3] has been pre-trained on 20,000 paired chest CT scans and reports. M3D [5] explores a versatile multi-modal large language model designed for universal 3D medical imaging analysis. fVLM [6] proposes a fine-grained pre-training framework that relies on segmentation to perform organ alignment. However, while these methods have demonstrated promising performance on various tasks such as zero-shot abnormality detection and report generation, several factors limit their training scalability and real-world applicability. For example, CT-CLIP and M3D are constrained to curated studies containing only a single imaging scan. Such studies rely heavily on human annotation and also fail to reflect real-world scenarios, where studies typically include multiple scans. fVLM further depends on a segmentation model, which introduces bias from segmentation quality and largely limits its scalability and generalizability. Moreover, these methods suffer from inefficient modeling of 3D medical imaging, leading to small batch sizes (e.g., 48), which are insufficient for effective language-image pre-training [2].
@@
Efficient Language-Image Pre-training is crucial, as it directly affects the number of sample pairs seen during training and contrasted per batch, two key factors of model capacity. Existing hierarchical architectures [12, 3] have been widely adopted to improve efficiency. However, as analyzed in Appendix D.2, these models may be less efficient than commonly assumed in the context of 3D medical imaging, even when compared with the original ViT. From the architecture perspective, components such as relative position embedding are expensive for 3D inputs and is not compatible with recent advancements like flash attention [18]. From the training perspective, FLIP [19] demonstrates a favorable trade-off by randomly removing 50% of tokens during training. However, models that require a fixed activation shape [12, 17, 38, 13] cannot benefit from this strategy. Moreover, many works in radiology [39, 9, 40, 6] have demonstrated the benefits of the universal feature learned by MAE [41]. Therefore, modeling 3D medical imaging with minimal adaptation of the original ViT [11] is a promising research direction and stands to benefit from recent advancements such as flash attention, patch dropout, and pre-trained models.
@@ 3 Method
Our goal is to perform language-image pre-training on uncurated studies. Given a study $S\in\mathbb{R}^{M\times 1\times D\times H\times W}$, where $M$ denotes the number of single-channel 3D scans, and $D$, $H$, and $W$ represent the depth, height, and width dimensions of each scan, respectively. Following the ViT [11], each scan is divided into a grid of non-overlapping volumes of size $(\frac{D}{d},\frac{H}{h},\frac{W}{w})$. These volumes are projected into visual tokens $F_{s}\in\mathbb{R}^{N\times c}$, where $N=M\times d\times h\times w$ is the total number of tokens and $c$ denotes the number of channels. As $N$ can be on the order of $10^{4}$, computing self-attention over all tokens throughout the ViT backbone is prohibitive in memory and also limits performance [15]. We introduce a novel hierarchical attention mechanism guided by the inherent data hierarchy of uncurated studies, enabling lighter self-attention while introducing effective priors derived from this hierarchy: slice, scan, and study.
Figure 2: Illustration of (a) the radiology data hierarchy for a single patient, including the study, single scan, and adjacent slices. Our hierarchical attention mechanism mirrors this hierarchy and computes self-attention independently within each level. (b) Our HLIP framework incorporates a visual encoder that performs attention at different levels. In practice, lightweight slice or scan attention with a few study attention layers suffices to extract features from the full study.
@@ 3.1 Hierarchical Attention Mechanism
The radiology data exhibits an inherent hierarchical structure for each patient, and our proposed hierarchical attention mechanism mirrors this structure. As illustrated in Figure 2, the data for an individual patient comprises three levels:
@@
  • The scan contains $D$ slices. While a single scan conveys only partial context of the full study, it still captures the complete extent of the target pathology.
@@
  • Study attention computes a single self-attention operation over all $N$ tokens in a study. The I/O complexity [42] of the study attention is $\Omega(N^{2}+N\times c)$ [18], as detailed in Appendix D.
@@
The proposed hierarchical attention mechanism contrasts with existing methods that modify attention computation using convolution [16, 17], pooling [13], or shifted windows [12], all of which require either regular activation shapes or attention masks. Our mechanism relies solely on a simple reshape operation, as there is no overlap between different scans or groups of adjacent slices [^1] With the original attention layer unchanged, study attention takes the input of size $(B,\ M\times d\times h\times w,\ c)$; scan attention takes the input of size $(B\times M,\ d\times h\times w,\ c)$; and slice attention takes the input of size $(B\times M\times d,\ h\times w,\ c)$, where $B$ denotes the batch size., allowing greater flexibility for uncurated studies that comprise multiple scans. As shown in Figure 2, each layer in the ViT can perform attention at any level. In practice, we evenly divide the ViT backbone into four subsets of layers (e.g., three layers per subset for the 12-layer ViT-B) and apply study attention only to the last layer of each subset, while the remaining layers perform the lighter scan or slice attention. Moreover, this intentionally simple design also enables seamless integration with recent efficiency advances such as patch dropout [19] and flash attention [18].